# Multiview learning with twin parametric margin SVM

### v1

@@ Abstract
Multiview learning (MVL) seeks to leverage the benefits of diverse perspectives to complement each other, effectively extracting and utilizing the latent information within the dataset. Several twin support vector machine-based MVL (MvTSVM) models have been introduced and demonstrated outstanding performance in various learning tasks. However, MvTSVM-based models face significant challenges in the form of computational complexity due to four matrix inversions, the need to reformulate optimization problems in order to employ kernel-generated surfaces for handling non-linear cases, and the constraint of uniform noise assumption in the training data. Particularly in cases where the data possesses a heteroscedastic error structure, these challenges become even more pronounced. In view of the aforementioned challenges, we propose multiview twin parametric margin support vector machine (MvTPMSVM). MvTPMSVM constructs parametric hyperplanes with the goal of maximizing the parametric margin between two classes, aiming to regulate and manage the impact of the heteroscedastic noise structure existing within the data. The proposed MvTPMSVM model avoids the explicit computation of matrix inversions in the dual formulation, leading to enhanced computational efficiency. We perform an extensive assessment of the MvTPMSVM model using benchmark datasets such as UCI, KEEL, synthetic, and Animals with Attributes (AwA). Our experimental results, coupled with rigorous statistical analyses, confirm the superior generalization capabilities of the proposed MvTPMSVM model compared to the baseline models. The source code of the proposed MvTPMSVM model is available at https://github.com/mtanveer1/MvTPMSVM.
@@ 1 Introduction
While the TSVM and its variants assume that the level of noise in the training data is consistent across the entire dataset [21]. Noise refers to random or irrelevant data that can interfere with the learning process by introducing errors or misleading patterns in the training data. The assumption of a uniform noise model may not hold true in real-world scenarios. In the case of a heteroscedastic noise structure, where noise strongly depends on the input value, the amount of noise varies based on the location. To generate a parametric-margin model, Hao [9] proposed the parametric margin $\nu$-support vector machine (par-$\nu$-SVM). This can be valuable in numerous scenarios, particularly when dealing with datasets exhibiting a heteroscedastic error structure, where the input values are greatly influenced by the noise. However, the par-$\nu$-SVMs learning rate is as slow as the standard SVM. To reduce the computational complexity, Peng [17] proposed twin parametric-margin SVM (TPMSVM). TPMSVM constructs two distinct nonparallel hyperplanes, with each of them serving as the parametric-margin hyperplane for the negative and positive classes.
@@
The efficacy of MVL algorithms hinges on two fundamental principles: the consensus principle and the complementarity principle. These principles are essential in determining how multiple views are related and combined. The complementarity principle highlights the importance of leveraging the complementary information provided by different views to ensure a comprehensive description of the data. The consensus principle aims to maximize the correlation among multiple distinct perspectives to achieve accurate classifiers for each individual view. In the domain of MVL, these two principles play a pivotal role in guiding the construction of models. Multiview support vector machine (SVM-2K) [6] is a consensus-based principle. SVM-2K model involves training two SVMs within distinct feature spaces, employing an additional $l_{1}$-norm constraint during the training process to uphold robustness. Considering the Rademacher complexity, incorporating the $l_{1}$-norm constraint results in a more rigorous generalization error bound for SVM-2K [6]. The Multiview twin support vector machine (MvTSVM) [41] is inspired by the co-regularization concept found in SVM-2K and utilizes the effective training attributes derived from TSVMs. By solving two smaller-sized QPPs, MvTSVM generates four non-parallel hyperplanes (associated with two views) in total, which makes it four times faster than SVM-2K. In recent years, several variants of MvTSVM have been proposed such as multiview twin hypersphere support vector machine (MvTHSVM) [48], multiview learning model based on nonparallel SVM (MvNPSVM) [28], multiview robust double-sided TSVM (MvRDTSVM) [46], deep multiview TSVM (DMvTSVM) [42], multiview one-class SVM method with privileged information learning (MOCPIL) [39], multiview large margin distribution machine (MVLDM) [10] and many more. However, MvTSVM encounters significant challenges, which are as follows: (1) MvTSVM necessitates four matrix inversion operations during the training process, resulting in avoidable computational overhead. (2) MvTSVM requires the reformulation of optimization problems when the training data is mapped into high-dimensional spaces using non-linear functions. (3) The uniformity of noise in the training data, or its pre-known functional dependency, presents significant obstacles to the effectiveness of the model.
@@
    We propose a novel multiview twin parametric margin support vector machine (MvTPMSVM). The proposed MvTPMSVM generates parametric hyperplanes that maximize the margin between the two classes, thereby controlling the impact of heteroscedastic noise present in the data.
@@ 2 Related Work
In this section, we briefly outline the mathematical formulations of SVM-2K and MvTSVM. Consider the sample space denoted as $\mathscr{T}$, which is a product of two distinct feature views, $A$ and $B$, expressed as $\mathscr{T}=\mathscr{T}^{A}\times\mathscr{T}^{B}$, where, $\mathscr{T}^{A}\in\mathbb{R}^{m_{1}}$ $\mathscr{T}^{B}\in\mathbb{R}^{m_{2}}$ and the label space is presented by $\mathscr{Y}\in\{-1,+1\}$. Suppose =$\mathscr{H}=\{(x_{i}^{A},x_{i}^{B},y_{i})|x_{i}^{A}\in\mathscr{T}^{A},x_{i}^{B}\in\mathscr{T}^{B},y_{i}\in\mathscr{Y}\}_{i=1}^{m}$ represent a two-view data set. Let us consider the input matrices $\textbf{A}^{A}$ and $\textbf{B}^{A}$ represent the $+1$ and $-1$ class sample of view $A$, and $\textbf{A}^{B}$ and $\textbf{B}^{B}$ represent the $+1$ and $-1$ class sample of view $B$, respectively. The non-parallel hyperplanes are given by $h_{1}^{A}=<w_{1}^{A},x^{A}>+b_{1}^{A}$ and $h_{2}^{A}=<w_{2}^{A},x^{A}>+b_{2}^{A}$ for view $A$, and two non-parallel hyperplanes for view $B$ is given by $h_{1}^{B}=<w_{1}^{B},x^{B}>+b_{1}^{B}$ and $h_{2}^{B}=<w_{2}^{B},x^{B}>+b_{2}^{B}$. Let $A_{1}=[\textbf{A}^{A},\hskip 2.84544pte_{1}]$, $B_{1}=[\textbf{A}^{B},\hskip 2.84544pte_{1}]$, $A_{2}=[\textbf{B}^{A},\hskip 2.84544pte_{2}]$, $B_{2}=[\textbf{B}^{B},\hskip 2.84544pte_{2}]$, $v_{1}=[w_{1}^{A};\hskip 2.84544ptb_{1}^{A}]$, $v_{2}=[w_{1}^{B};\hskip 2.84544ptb_{1}^{B}]$, $u_{1}=[w_{2}^{A};\hskip 2.84544ptb_{2}^{A}]$, and $u_{2}=[w_{2}^{B};\hskip 2.84544ptb_{2}^{B}]$, where $e_{1}\in\mathbb{R}^{m_{1}}$, $e_{2}\in\mathbb{R}^{m_{2}}$ are the vector of ones, respectively. Table 1 provides a summary of the primary notations used in this paper.
@@ 2.1 Two view learning: SVM-2K, theory and practice
s.t. $\lvert<w^{B},\phi(x_{i}^{B})>+b^{B}-<w^{A},\phi(x_{i}^{A})>-b^{A}\rvert\leq\epsilon+\eta_{i},$
$y_{i}(<w^{B},\phi(x_{i}^{B})>+b^{B})\geq 1-\xi_{i}^{B},$
$y_{i}(<w^{A},\phi(x_{i}^{A})>+b^{A})\geq 1-\xi_{i}^{A},$
@@
where $C_{1}$, $C_{2}$, $D$ $(>0)$ are penalty parameters, $\epsilon>0$ is an insensitive parameter and $\xi_{i}^{A}$, $\xi_{i}^{B}$, and $\eta_{i}$ are slack variables. The amalgamation of the two views through the $\epsilon$-insensitive $l_{1}$-norm similarity constraint is incorporated as the initial constraint in the problem (2.1).
The dual problem of (2.1) is given by:
[…]
s.t. $\xi_{i}^{A}=\alpha_{i}^{A}y_{i}-\beta_{i}^{+}+\beta_{i}^{-},$
$\xi_{i}^{B}=\alpha_{i}^{B}y_{i}+\beta_{i}^{+}-\beta_{i}^{-},$
$\sum_{i=1}^{n}\xi_{i}^{A}=\sum_{i=1}^{n}\xi_{i}^{B}=0,$
@@ 2.2 Multiview twin support vector machine (MvTSVM)
MvTSVM [41] needs to construct two non-parallel hyperplanes. The optimization problem of MvTSVM is given as follows:
@@
where $C_{1}$, $C_{2}$, $C_{3}$, $C_{4}$, $D_{1}$, $D_{2}$ and $\epsilon$ $(>0)$ are tunable parameters; $\xi_{1}$, $\xi_{2}$, $\xi_{3}$, and $\xi_{4}$ are slack variables respectively.
@@
$\underset{\xi_{1},\leavevmode\nobreak\ \xi_{2},\leavevmode\nobreak\ \alpha_{1},\leavevmode\nobreak\ \alpha_{2}}{\min}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\xi_{1}^{T}(A_{1}^{T}A_{1})^{-1}\xi_{1}+\xi_{2}^{T}(A_{2}^{T}A_{2})^{-1}\xi_{2}-(\alpha_{1}+\alpha_{2})^{T}e_{2}$
s.t. $\xi_{1}=A_{1}^{T}(\beta_{2}-\beta_{1})-B_{1}^{T}\alpha_{1},$
$\xi_{2}=A_{2}^{T}(\beta_{1}-\beta_{2})-B_{2}^{T}\alpha_{2},$
@@
$\underset{\eta_{1},\leavevmode\nobreak\ \eta_{2},\leavevmode\nobreak\ \gamma_{1},\leavevmode\nobreak\ \gamma_{2}}{\min}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\eta_{1}^{T}(B_{1}^{T}B_{1})^{-1}\eta_{1}+\eta_{2}^{T}(B_{2}^{T}B_{2})^{-1}\eta_{2}-(\gamma_{1}+\gamma_{2})^{T}e_{1}$
s.t. $\eta_{1}=B_{1}^{T}(\delta_{2}-\delta_{1})-A_{1}^{T}\gamma_{1},$
$\eta_{2}=B_{2}^{T}(\delta_{1}-\delta_{2})-A_{2}^{T}\gamma_{2},$
@@ 3.1 Linear MvTPMSVM
    MvTPMSVM operates as a data projection model that translates information from the feature space into a one-dimensional space, effectively representing class labels. By leveraging label correlations, it incorporates the $\epsilon$-insensitive loss function, which gives rise to the inclusion of a regularization term between different views. Nonnegative slack variables $\eta_{1}$ and $\eta_{2}$ are introduced, along with tunable parameters $D_{1}$ and $D_{2}$, to evaluate the degree of deviation from $\epsilon$-similarity between two positive (or negative) classifiers across two views.
@@
    The first constraint in the objective function of the proposed MvTPMSVM model is the $\epsilon$-insensitive $l_{1}$-norm, which ensures the alignment of the classifiers from both views concerning positive and negative classes. Samples deviating from these constraints are accommodated by introducing the parameter $\epsilon$. By leveraging label correlations, this approach diminishes the gap between different perspectives, thereby ensuring their consistency.
@@
    The second component of the objective function is made to minimize the total projection values of the positive (negative) training points in each view. Optimizing this component encourages the effective positioning of negative (positive) training points to attain maximal distance from the positive (negative) parametric-margin hyperplane.
@@
    The second and third constraints in the optimization problem of the proposed MvTPMSVM model stipulate that the parametric-margin hyperplane should not be less than zero. The fault tolerance parameters $\xi_{1}$, $\xi_{2}$, $\xi_{3}$, and $\xi_{4}$ are employed to quantify the extent of disagreement on each view. In order to avoid overfitting to the positive (negative) training points, the third term of the objective function minimizes the sum of error variables.
@@
$\underset{\alpha_{1}^{+},\leavevmode\nobreak\ \alpha_{2}^{+},\leavevmode\nobreak\ \beta_{1}^{+},\leavevmode\nobreak\ \beta_{2}^{+}}{\min}$ $-\frac{1}{2}(\alpha_{1}^{+}-\beta_{1}^{+}+\beta_{2}^{+})^{T}A_{1}A_{1}^{T}(\alpha_{1}^{+}-\beta_{1}^{+}+\beta_{2}^{+})$
@@
$\underset{\alpha_{1}^{-},\leavevmode\nobreak\ \alpha_{2}^{-},\leavevmode\nobreak\ \beta_{1}^{-},\leavevmode\nobreak\ \beta_{2}^{-}}{\min}$ $-\frac{1}{2}(\alpha_{1}^{-}-\beta_{1}^{-}+\beta_{2}^{-})^{T}B_{1}B_{1}^{T}(\alpha_{1}^{-}-\beta_{1}^{-}+\beta_{2}^{-})$
@@
$\underset{\tilde{\tau}}{\min}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\tilde{\tau}^{T}\tilde{\Lambda}\tilde{\tau}+\kappa_{1}\tilde{\tau}$
@@
$\underset{\hat{\tau}}{\min}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\hat{\tau}^{T}\hat{\Lambda}\hat{\tau}+\kappa_{2}\hat{\tau}$

### v2

@@ Abstract
Multiview learning (MVL) seeks to leverage the benefits of diverse perspectives to complement each other, effectively extracting and utilizing the latent information within the dataset. Several twin support vector machine-based MVL (MvTSVM) models have been introduced and demonstrated outstanding performance in various learning tasks. However, MvTSVM-based models face significant challenges in the form of computational complexity due to four matrix inversions, the need to reformulate optimization problems in order to employ kernel-generated surfaces for handling non-linear cases, and the constraint of uniform noise assumption in the training data. Particularly in cases where the data possesses a heteroscedastic error structure, these challenges become even more pronounced. In view of the aforementioned challenges, we propose multiview twin parametric margin support vector machine (MvTPMSVM). MvTPMSVM constructs parametric margin hyperplanes corresponding to both classes, aiming to regulate and manage the impact of the heteroscedastic noise structure existing within the data. The proposed MvTPMSVM model avoids the explicit computation of matrix inversions in the dual formulation, leading to enhanced computational efficiency. We perform an extensive assessment of the MvTPMSVM model using benchmark datasets such as UCI, KEEL, synthetic, and Animals with Attributes (AwA). Our experimental results, coupled with rigorous statistical analyses, confirm the superior generalization capabilities of the proposed MvTPMSVM model compared to the baseline models. The source code of the proposed MvTPMSVM model is available at https://github.com/mtanveer1/MvTPMSVM.
@@ 1 Introduction
While the TSVM and its variants assume that the level of noise in the training data is consistent across the entire dataset [21]. Noise refers to random or irrelevant data that can interfere with the learning process by introducing errors or misleading patterns in the training data. The assumption of a uniform noise model may not hold true in real-world scenarios. In the case of a heteroscedastic noise structure, where noise strongly depends on the input value, the amount of noise varies based on the location. To generate a parametric-margin model, Hao [9] proposed the parametric margin $\nu$-support vector machine (par-$\nu$-SVM). This can be valuable in numerous scenarios, particularly when dealing with datasets exhibiting a heteroscedastic error structure, where the input values are greatly influenced by the noise. However, the par-$\nu$-SVMs learning rate is as slow as the standard SVM. To reduce the computational complexity, Peng [17] proposed twin parametric-margin SVM (TPMSVM). TPMSVM constructs two distinct nonparallel hyperplanes, with each of them serving as the parametric-margin hyperplane for the $+1$ and $-1$ classes.
@@
The efficacy of MVL algorithms hinges on two fundamental principles: the consensus principle and the complementarity principle. These principles are essential in determining how multiple views are related and combined. The complementarity principle highlights the importance of leveraging the complementary information provided by different views to ensure a comprehensive description of the data. The consensus principle aims to maximize the correlation among multiple distinct perspectives to achieve accurate classifiers for each individual view. In the domain of MVL, these two principles play a pivotal role in guiding the construction of models. Multiview SVM (SVM-2K) [6] is a consensus-based principle. SVM-2K model involves training two SVMs within distinct feature spaces, employing an additional $l_{1}$-norm constraint during the training process to uphold robustness. Considering the Rademacher complexity, incorporating the $l_{1}$-norm constraint results in a more rigorous generalization error bound for SVM-2K [6]. The Multiview TSVM (MvTSVM) [41] is inspired by the co-regularization concept found in SVM-2K and utilizes the effective training attributes derived from TSVMs. By solving two smaller-sized QPPs, MvTSVM generates four non-parallel hyperplanes (associated with two views) in total, which makes it four times faster than SVM-2K. In recent years, several variants of MvTSVM have been proposed such as multiview twin hypersphere support vector machine (MvTHSVM) [48], multiview learning model based on nonparallel SVM (MvNPSVM) [28], multiview robust double-sided TSVM (MvRDTSVM) [46], deep multiview TSVM (DMvTSVM) [42], multiview one-class SVM method with privileged information learning (MOCPIL) [39], multiview large margin distribution machine (MVLDM) [10] and many more. However, MvTSVM encounters significant challenges, which are as follows: (1) MvTSVM necessitates four matrix inversion operations during the training process, resulting in avoidable computational overhead. (2) MvTSVM requires the reformulation of optimization problems when the training data is mapped into high-dimensional spaces using non-linear functions. (3) The uniformity of noise in the training data, or its pre-known functional dependency, presents significant obstacles to the effectiveness of the model.
@@
    We propose a novel multiview twin parametric margin support vector machine (MvTPMSVM). The proposed MvTPMSVM generates parametric margin hyperplanes corresponding to both classes, thereby controlling the impact of heteroscedastic noise present in the data.
@@ 2 Related Work
In this section, we briefly outline the mathematical formulations of SVM-2K and MvTSVM. Consider the sample space denoted as $\mathscr{T}$, which is a product of two distinct feature views, $A$ and $B$, expressed as $\mathscr{T}=\mathscr{T}^{A}\times\mathscr{T}^{B}$, where, $\mathscr{T}^{A}\subseteq\mathbb{R}^{m_{1}}$ $\mathscr{T}^{B}\subseteq\mathbb{R}^{m_{2}}$ and the label space is presented by $\mathscr{Y}=\{-1,+1\}$. Suppose =$\mathscr{H}=\{(x_{i}^{A},x_{i}^{B},y_{i})|x_{i}^{A}\in\mathscr{T}^{A},x_{i}^{B}\in\mathscr{T}^{B},y_{i}\in\mathscr{Y}\}_{i=1}^{m}$ represent a two-view data set. Let us consider the input matrices $\textbf{A}^{A}$ and $\textbf{B}^{A}$ represent the $+1$ and $-1$ class sample of view $A$, and $\textbf{A}^{B}$ and $\textbf{B}^{B}$ represent the $+1$ and $-1$ class sample of view $B$, respectively. The non-parallel hyperplanes are given by $h_{1}^{A}=<w_{1}^{A},x^{A}>+b_{1}^{A}$ and $h_{2}^{A}=<w_{2}^{A},x^{A}>+b_{2}^{A}$ for view $A$, and two non-parallel hyperplanes for view $B$ is given by $h_{1}^{B}=<w_{1}^{B},x^{B}>+b_{1}^{B}$ and $h_{2}^{B}=<w_{2}^{B},x^{B}>+b_{2}^{B}$. Let $A_{1}=[\textbf{A}^{A},\hskip 2.84544pte_{1}]$, $B_{1}=[\textbf{A}^{B},\hskip 2.84544pte_{1}]$, $A_{2}=[\textbf{B}^{A},\hskip 2.84544pte_{2}]$, $B_{2}=[\textbf{B}^{B},\hskip 2.84544pte_{2}]$, $v_{1}=[w_{1}^{A};\hskip 2.84544ptb_{1}^{A}]$, $v_{2}=[w_{1}^{B};\hskip 2.84544ptb_{1}^{B}]$, $u_{1}=[w_{2}^{A};\hskip 2.84544ptb_{2}^{A}]$, and $u_{2}=[w_{2}^{B};\hskip 2.84544ptb_{2}^{B}]$, where $e_{1}\in\mathbb{R}^{m_{1}}$, $e_{2}\in\mathbb{R}^{m_{2}}$ are the vector of ones, respectively. Table 1 provides a summary of the primary notations used in this paper.
@@ 2.1 Two view learning: SVM-2K, theory and practice
s.t. $\lvert<w^{B},\phi_{B}(x_{i}^{B})>+b^{B}-<w^{A},\phi_{A}(x_{i}^{A})>-b^{A}\rvert\leq\epsilon+\eta_{i},$
$y_{i}(<w^{B},\phi_{B}(x_{i}^{B})>+b^{B})\geq 1-\xi_{i}^{B},$
$y_{i}(<w^{A},\phi_{A}(x_{i}^{A})>+b^{A})\geq 1-\xi_{i}^{A},$
@@
where $C_{1}$, $C_{2}$, $D$ $(>0)$ are penalty parameters, $\epsilon>0$ is an insensitive parameter and $\xi_{i}^{A}$, $\xi_{i}^{B}$, and $\eta_{i}$ are slack variables, respectively. The amalgamation of the two views through the $\epsilon$-insensitive $l_{1}$-norm similarity constraint is incorporated as the initial constraint in the problem (2.1). The dual problem of (2.1) is given by:
$\underset{\beta_{i}^{+},\leavevmode\nobreak\ \beta_{i}^{-},\leavevmode\nobreak\ \alpha_{i}^{A},\leavevmode\nobreak\ \alpha_{i}^{B}}{\max}\hskip 1.42271pt\leavevmode\nobreak$ $-\frac{1}{2}\sum_{i,j=1}^{m}(g_{i}^{A}g_{j}^{A}k_{A}(x_{i},x_{j})+g_{i}^{B}g_{j}^{B}k_{B}(x_{i},x_{j}))+\sum_{i=1}^{m}(\alpha_{i}^{A}+\alpha_{i}^{B})$
s.t. $g_{i}^{A}=\alpha_{i}^{A}y_{i}-\beta_{i}^{+}+\beta_{i}^{-},$
$g_{i}^{B}=\alpha_{i}^{B}y_{i}+\beta_{i}^{+}-\beta_{i}^{-},$
$\sum_{i=1}^{n}g_{i}^{A}=\sum_{i=1}^{n}g_{i}^{B}=0,$
@@ 2.2 Multiview twin support vector machine (MvTSVM)
MvTSVM [41] needs to construct four non-parallel hyperplanes. The optimization problem of MvTSVM is given as follows:
@@
where $C_{1}$, $C_{2}$, $C_{3}$, $C_{4}$, $D_{1}$, $D_{2}$ are tunable parameters, and $\epsilon>0$ is an insensitive parameter; $\xi_{1}$, $\xi_{2}$, $\xi_{3}$, $\xi_{4}$, $\eta_{1}$, and $\eta_{2}$ are slack variables, respectively.
@@
$\underset{\beta_{1},\leavevmode\nobreak\ \beta_{2},\leavevmode\nobreak\ \alpha_{1},\leavevmode\nobreak\ \alpha_{2}}{\max}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\zeta_{1}^{T}(A_{1}^{T}A_{1})^{-1}\zeta_{1}+\zeta_{2}^{T}(A_{2}^{T}A_{2})^{-1}\zeta_{2}-(\alpha_{1}+\alpha_{2})^{T}e_{2}$
s.t. $\zeta_{1}=A_{1}^{T}(\beta_{2}-\beta_{1})-B_{1}^{T}\alpha_{1},$
$\zeta_{2}=A_{2}^{T}(\beta_{1}-\beta_{2})-B_{2}^{T}\alpha_{2},$
@@
$\underset{\delta_{1},\leavevmode\nobreak\ \delta_{2},\leavevmode\nobreak\ \gamma_{1},\leavevmode\nobreak\ \gamma_{2}}{\max}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\rho_{1}^{T}(B_{1}^{T}B_{1})^{-1}\rho_{1}+\rho_{2}^{T}(B_{2}^{T}B_{2})^{-1}\rho_{2}-(\gamma_{1}+\gamma_{2})^{T}e_{1}$
s.t. $\rho_{1}=B_{1}^{T}(\delta_{2}-\delta_{1})-A_{1}^{T}\gamma_{1},$
$\rho_{2}=B_{2}^{T}(\delta_{1}-\delta_{2})-A_{2}^{T}\gamma_{2},$
@@ 3.1 Linear MvTPMSVM
    The second term in the objective function minimizes the sum of projection values for the $-1$ class training samples of both views. By leveraging label correlations, it incorporates the $\epsilon$-insensitive loss function, which gives rise to the inclusion of a regularization term between different views. Nonnegative slack variables $\eta_{1}$ and $\eta_{2}$ are introduced, along with tunable parameters $D_{1}$ and $D_{2}$, to evaluate the degree of deviation from $\epsilon$-similarity between two $+1$ ($-1$) classifiers across two views.
@@
    The first constraint in the objective function of the proposed MvTPMSVM model is the $\epsilon$-insensitive $l_{1}$-norm, which ensures the alignment of the classifiers from both views concerning $+1$ and $-1$ classes. Samples deviating from these constraints are accommodated by introducing the parameter $\epsilon$. By leveraging label correlations, this approach diminishes the gap between different perspectives, thereby ensuring their consistency.
@@
    The second component of the objective function is made to minimize the total projection values of the $+1$ ($-1$) training points in each view. Optimizing this component encourages the effective positioning of $-1$ ($+1$) training points to attain maximal distance from the $+1$ ($-1$) parametric-margin hyperplane.
@@
    The second and third constraints in the optimization problem of the proposed MvTPMSVM model stipulate that the parametric-margin hyperplane should not be less than zero. The fault tolerance parameters $\xi_{1}$, $\xi_{2}$, $\xi_{3}$, and $\xi_{4}$ are employed to quantify the extent of disagreement on each view. In order to avoid overfitting to the $+1$ ($-1$) training points, the third term of the objective function minimizes the sum of error variables.
@@
$\underset{\alpha_{1}^{+},\leavevmode\nobreak\ \alpha_{2}^{+},\leavevmode\nobreak\ \beta_{1}^{+},\leavevmode\nobreak\ \beta_{2}^{+}}{\max}$ $-\frac{1}{2}(\alpha_{1}^{+}-\beta_{1}^{+}+\beta_{2}^{+})^{T}A_{1}A_{1}^{T}(\alpha_{1}^{+}-\beta_{1}^{+}+\beta_{2}^{+})$
@@
$\underset{\alpha_{1}^{-},\leavevmode\nobreak\ \alpha_{2}^{-},\leavevmode\nobreak\ \beta_{1}^{-},\leavevmode\nobreak\ \beta_{2}^{-}}{\max}$ $-\frac{1}{2}(\alpha_{1}^{-}-\beta_{1}^{-}+\beta_{2}^{-})^{T}B_{1}B_{1}^{T}(\alpha_{1}^{-}-\beta_{1}^{-}+\beta_{2}^{-})$
@@
$\underset{\tilde{\tau}}{\max}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\tilde{\tau}^{T}\tilde{\Lambda}\tilde{\tau}+\kappa_{1}\tilde{\tau}$
@@
$\underset{\hat{\tau}}{\max}\hskip 2.84544pt\leavevmode\nobreak$ $\frac{1}{2}\hat{\tau}^{T}\hat{\Lambda}\hat{\tau}+\kappa_{2}\hat{\tau}$